OpenStudy (anonymous):

refer to the expressions x+75, x+5, x-10, x, x+100 a) arrange the data from smallest to largest b)what is x if the median is 200

OpenStudy (dumbcow):

x-10,x, x+5, x+75, x+100 median is middle number median = x+5 = 200 x=195

OpenStudy (anonymous):

how do you know that x+5 is the median

OpenStudy (dumbcow):

because its the middle number

OpenStudy (anonymous):

the middle number of what?

OpenStudy (dumbcow):

the list of numbers if you have the list of {1,2,3} 2 is the middle value, greater than1 and less than 3

OpenStudy (anonymous):

okay i get that but where's the list that you used to find out that x+5 is the median

OpenStudy (dumbcow):

you cant remember the list of numbers you just posted?

OpenStudy (anonymous):

x+75, x+5, x-10, x, x+100 okay but i get x-10 as the median not x+5. can you explain this?

OpenStudy (dumbcow):

i can, place them in order from least to greatest first which i already showed you earlier x-10,x, x+5, x+75, x+100

OpenStudy (anonymous):

oh okay. i get it now. so that would be the answer to part a. now whats the answer to part b

OpenStudy (dumbcow):

median is middle number median = x+5 = 200 x=195

OpenStudy (anonymous):

how do you get 195?

OpenStudy (dumbcow):

oh im sorry i forgot you were an idiot 200 - 5 = 195

OpenStudy (anonymous):

jackass. where did the x go?

OpenStudy (anonymous):

i'm really hungry

OpenStudy (dumbcow):

somewhere between w and y

OpenStudy (anonymous):

wow i didn't know cows had a sense of humor

OpenStudy (dumbcow):

go integrate yourself

OpenStudy (anonymous):

well i think im gonna just derive myself outta here

OpenStudy (anonymous):

hey what do you want to eat for dinner?

OpenStudy (dumbcow):

cat

OpenStudy (anonymous):

i was thinking more like beef

OpenStudy (dumbcow):

please leave...NOW